### v1

# GARCH copulas and GARCH-mimicking copulas

Abstract
The bivariate copulas that describe the dependencies and partial dependencies of lagged variables in strictly stationary, first-order GARCH-type processes are investigated. It is shown that the copulas of symmetric GARCH processes are jointly symmetric but non-exchangeable, while the copulas of processes with symmetric innovation distributions and asymmetric leverage effects have weaker h-symmetry; copulas with asymmetric innovation distributions have neither form of symmetry. Since the actual copulas are typically inaccessible, due to the unknown functional forms of the marginal distributions of GARCH processes, methods of mimicking them are proposed. These rely on constructions that combine standard bivariate copulas for positive dependence with two uniformity-preserving transformations known as v-transforms. A variety of new copulas are introduced and the ones providing the best fit to simulated data from GARCH-type processes are identified. A method of constructing tractable simplified d-vines using linear v-transforms is described and shown to coincide with the vt-d-vine model when the two v-transforms are identical.
Keywords: Time series; copulas; vine copulas; GARCH processes; v-transforms.
1 Introduction
The ARCH process (Engle,, 1982), the GARCH process (Bollerslev,, 1986) and a whole lexicon of related GARCH-type models have proved to be remarkably successful models for time series displaying stochastic volatility, such as asset price return series. In this paper we analyse the serial dependence structure of first-order GARCH-type models in terms of the implied bivariate copulas that describe the dependence of pairs of variables at different lags. Our aim is to understand how we could mimic such dependence structures with appropriately chosen bivariate copulas arranged in d-vines; we do this with a view to developing models that might offer additional flexibility in practice. The approach we take recognizes that uniformity-preserving transformations, known as v-transforms (McNeil,, 2021), are embedded in the implied copulas of classical GARCH models and suggests that they could be used as explicit modelling elements to construct vines for applications.
The essential mechanism of a GARCH-type model is simple and intuitive; the conditional distribution from which a new observation is drawn has a standard deviation that is a function of the size of previous returns, so that large market moves have a tendency to beget further large market moves and create periods of volatility. The ARCH and GARCH prototypes of Engle, (1982) and Bollerslev, (1986) have inspired a vast class of related GARCH-type models, including the GJR-GARCH model, which incorporates leverage effects to reflect different contributions to volatility by returns of differing signs (Glosten et al.,, 1993), power-GARCH models (Ding et al.,, 1993) and exponential GARCH models (Nelson,, 1991), to name a but a few of the more important extensions. A number of comprehensive surveys of the extended GARCH family have been published including Bollerslev et al., (1992) Bollerslev et al., (1994) and Shephard, (1996).
Notwithstanding the appealing basic mechanism for explaining volatility, GARCH-type models are also something of a black box. The stationary distribution is not generally known in closed form, although it is known that many models gives rise to distributions with Pareto tails (Mikosch and Stărică,, 2000). In particular, the serial dependence structure is not well understood. If $(X_{t})_{t\in\mathbb{Z}}$ follows a stationary GARCH-type model then the distribution of $(X_{1},\ldots,X_{n})^{\top}$ for $n\geqslant 2$ must have a copula, but this copula also has no simple, accessible closed form.
It may be asked why we would want to mimic GARCH-type models with vine copulas, given that GARCH models have proved themselves in practice over four decades. While many asset return series are well modelled by GARCH, we do encounter series where a power-lawed marginal distribution is sub-optimal for the data; see, for example, Bladt and McNeil, 2022b . For such datasets it would be convenient to be able to construct models that combined bespoke choices of marginal distribution, or non-parametric margins, with a copula that could mimic the kind of serial dependence that is implicit in the GARCH family. If we understand the types of copula that are implicit in the GARCH family, it may also be possible to improve upon these copulas by adding further flexibility.
Vine copulas also have a large literature, including Joe, (1996, 1997), Bedford and Cooke, (2001, 2002), Kurowicka and Cooke, (2006), Aas et al., (2009) and Smith et al., (2010). The most suitable vine structure for a time series is the d-vine, which is able to describe strict stationarity of a random vector under additional translation-invariance restrictions on the vine structure (Nagler et al.,, 2022). Markov copula models (Darsow et al.,, 1992; Chen and Fan,, 2006; Beare,, 2010; Domma et al.,, 2009) are simple examples of stationary first-order d-vines and a number of authors have written on their higher-order extensions (Ibragimov,, 2009; Beare and Seo,, 2015; Brechmann and Czado,, 2015; Loaiza-Maya et al.,, 2018; Nagler et al.,, 2022). The last-named paper introduces the terminology s-vine for stationary multivariate vines, which includes the stationary univariate d-vine as a special case.
The joint density of the distribution of $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary process, such as a stationary model of GARCH type, can be decomposed as a d-vine. In other words we may write
$f_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}f_{X}(x_{i})\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{X_{1}|X_{2},\ldots,X_{k}}(x_{j-k}\mid x_{j-k+1},\ldots,x_{j-1}),\\ F_{X_{k+1}|X_{2},\ldots,X_{k}}(x_{j}\mid x_{j-k+1},\ldots,x_{j-1})\mid x_{j-k+1},\ldots,x_{j-1})$ (1)
where $f_{X}$ denotes the marginal density, $F_{X_{i}|X_{2},\ldots,X_{k}}$ denotes the conditional distribution function of $X_{i}$ given $X_{2},\ldots,X_{k}$ and $c_{k}(u,v\mid x_{2},\ldots,x_{k})$ is the density of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $X_{2}=x_{2},\ldots,X_{k}=x_{k}$. In the case $k=1$ the conditional distribution is simply the marginal distribution $F_{X}$ and $c_{1}$ is simply the density of the copula of the unconditional distribution of $(X_{1},X_{2})^{\top}$. When the process is first-order Markov, all of the terms involving $c_{k}(u,v\mid\cdot)$ for $k>1$ disappear, but this will only apply to the ARCH(1)-type models among the models we analyse.
Although the densities of GARCH-type models may be written as d-vines, they may not in general be written as so-called simplified d-vines, that is models in which the conditional copula densities $c_{k}$ do not depend explicitly on the values $x_{j-k+1},\ldots,x_{j-1}$ in equation (1) for $k\geqslant 2$. When d-vines are used in practical modelling applications they are almost always simplified d-vines. However, as pointed out by a number of authors (Haff et al.,, 2010; Stöber et al.,, 2013; Spanhel and Kurz,, 2019; Mroz et al.,, 2021), these may not be able to capture all the features present in an arbitrary joint density, such as that of a GARCH-type model. In the course of our analysis we return to this question.
We confine our attention to first-order GARCH-type models. In practice models of order (1,1) often work very well and adding higher-order terms yields only minor improvements. We also look in some detail at ARCH(1)-type models for two reasons: they are first-order Markov models and therefore particularly easy to analyse; they give rise to the essential copula forms that are encountered in the more general class, as will be seen in this paper.
Loaiza-Maya et al., (2018) have previously investigated the copula of the classical ARCH(1) process. Since this model is first-order Markov, its serial dependence may be fully described by a first-order d-vine with a single bivariate copula, namely the copula of the distribution of two successive values $(X_{1},X_{2})^{\top}$. Using a simulation approach, they identify the cross-shaped nature of the copula density and suggest that mixtures of positive dependence and negative dependence copulas (such as a t copula and a 90-degree rotated t copula) can model the dependence structure effectively. Zhao et al., (2022) argue that t copulas in d-vines can approximate the dependence structure of GARCH(1,1) and GJR-GARCH(1,1) models. We will find some evidence in support of t copulas but also argue that they miss some important asymmetries.
We make a number of original contributions. First, we analyse the structure of the bivariate copula densities $c_{k}$ in (1) for classical GARCH-type models; we show that symmetric models give rise to jointly symmetric copulas while models with symmetric innovation distributions and asymmetric volatility have copulas with weaker symmetry, which we call h-symmetry. We confirm the theoretical insights by using numerical approximations to provide accurate contour plots of the copula densities. These reveal hitherto overlooked features, such as the non-exchangeability of the copulas. We also examine the effects of leverage, asymmetric innovations and autoregressive conditional mean terms on the copulas and on their symmetries.
Second, we show how uniformity-preserving transformations known as v-transforms (McNeil,, 2021) play an implicit role in the copulas of classic symmetric models and in models with certain forms of asymmetry. We investigate methods of constructing copulas with v-transforms to mimic the behaviour of the copulas implied by GARCH-type models. Using simulated data we identify the copulas based on v-transforms that provide the best fit to simulated data and show that these typically do better than mixture copulas.
Finally, we draw conclusions with respect to the construction of d-vine copulas to mimic the serial-dependence structure of GARCH-type models. We also show how d-vine models with appropriately chosen copulas can be related to the vt-ARMA and vt-d-vine models proposed by McNeil, (2021) and Bladt and McNeil, 2022b .
2 Preliminary results on symmetry
2.1 Definitions
Symmetry plays an important role in this paper and we give a definition to make our use of various symmetry concepts clear; parts (iii) and (iv) are from Nelsen, (2006).
Definition 1.
Let $(Y,Z)^{\top}$ be a generic random vector taking values in $\mathbb{R}^{2}$.
  1. (i)
    If $(Y,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,b-Z)^{\top}$ for some $b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is h-symmetric about $b$ (symmetric about a horizontal axis at $b$).
  2. (ii)
    If $(Y-a,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,Z)^{\top}$ for some $a\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is v-symmetric about $a$ (symmetric about a vertical axis at $a$).
  3. (iii)
    If $(Y-a,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,b-Z)^{\top}$ for $a,b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is radially symmetry about $(a,b)$.
  4. (iv)
    If $(Y,Z)^{\top}$ has all three of the above forms of symmetry then we say that it is jointly symmetric about $(a,b)$.
Remarks 1.
  1. 1.
    When $(Y,Z)^{\top}$ is h-symmetric about $b$, the marginal distribution of $Z$ is symmetric about $b$; when $(Y,Z)^{\top}$ is v-symmetric about $a$, the marginal distribution of $Y$ is symmetric about $a$. The reverse implications are not true.
  2. 2.
    It may be noted that we can always shift the distribution of $(Y,Z)^{\top}$ in two-dimensional space so that axes or centres of symmetry are at zero or the origin. Where $a$ and $b$ are not relevant to the result, we drop the explicit reference to them and just refer to h-symmetry, v-symmetry, radial symmetry and joint symmetry.
  3. 3.
    H-symmetry and v-symmetry together imply radial symmetry and therefore joint symmetry. However radial symmetry does not necessarily imply h-symmetry and v-symmetry, although it does imply symmetry of both marginal distributions.
The symmetries in Definition 1 have the following implications for the copula of $(Y,Z)^{\top}$.
Lemma 1.
Let $(Y,Z)^{\top}$ be a random vector with continuous marginal distributions and copula $C$.
  1. (i)
    If $(Y,Z)^{\top}$ is h-symmetric then $C(u,v)=u-C(u,1-v)$, for $(u,v)\in[0,1]^{2}$.
  2. (ii)
    If $(Y,Z)^{\top}$ is v-symmetric then $C(u,v)=v-C(1-u,v)$, for $(u,v)\in[0,1]^{2}$,
  3. (iii)
    If $(Y,Z)^{\top}$ is radially symmetric then $C(u,v)=u+v+C(1-u,1-v)-1$ , for $(u,v)\in[0,1]^{2}$.
Proof.
Since the copula is invariant under strictly increasing transformations of the margins, such as changes of location, we can, without loss of generality, assume that the axes and centres of symmetry are at zero and the origin. For case (i) observe that if $(Y,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,-Z)^{\top}$ then the joint distribution function satisfies
$F_{Y,Z}(y,z)=\mathbb{P}(Y\leqslant y,Z\leqslant z)=\mathbb{P}(Y\leqslant y,-Z\leqslant z)$ $=\mathbb{P}(Y\leqslant y)-P(Y\leqslant y,Z\leqslant-z)$
$=F_{Y}(y)-F_{Y,Z}(y,-z)$ (2)
for all $(y,z)\in\mathbb{R}^{2}$ and application of Sklar’s Theorem (Sklar,, 1959) implies that the (unique) copula $C$ must satisfy
$C(F_{Y}(y),F_{Z}(z))=F_{Y}(y)-C(F_{Y}(y),F_{Z}(-z))=F_{Y}(y)-C(F_{Y}(y),1-F_{Z}(z))$
where $F_{Y}$ and $F_{Z}$ denote the marginal distribution functions and the final equality follows from the symmetry of the marginal distribution of $Z$. Hence we obtain that $C(u,v)=u-C(u,1-v)$ at all points in the unit square. The proof of (ii) is obviously identical to (i) while (iii) is a well-known result found in Nelsen, (2006) (Theorem 2.7.3). ∎
Definition 2.
A copula $C$ satisfying $C(u,v)=u-C(u,1-v)$ for all $(u,v)\in[0,1]^{2}$ will be said to have h-symmetry. Similarly, a copula satisfying $C(u,v)=v-C(1-u,v)$ will be said to have v-symmetry. A copula satisfying $C(u,v)=u+v+C(1-u,1-v)-1$ will be said to have radial symmetry. A copula with all three forms of symmetry will be said to be jointly symmetric.
We note that joint symmetry of the copula is the bivariate special case of a property that Durante and Fuchs, (2019) call reflection invariance.
2.2 Correlation and rank correlation
Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined). Perhaps surprisingly, h-symmetry on its own is a sufficient condition for an absence of both correlation and rank correlation. The same is true of v-symmetry.
Proposition 1.
If a random vector $(Y,Z)^{\top}$ has finite second moments and is either h-symmetric or v-symmetric then the Pearson correlation $\rho(Y,Z)$ is zero.
Proof.
We give the proof in the case of h-symmetry, the method for v-symmetry being identical. If $(Y,Z)^{\top}$ is h-symmetric about $b$ then $\mathbb{E}(Z)=b$ and we can write $\operatorname{cov}(Y,Z)=\mathbb{E}(YZ)-\mathbb{E}(Y)\mathbb{E}(Z)=\mathbb{E}(Y(Z-b))+\mathbb{E}(Yb)-\mathbb{E}(Y)b=\mathbb{E}(Y(Z-b))$. However, h-symmetry implies $\mathbb{E}(Y(Z-b))=\mathbb{E}(Y(b-Z))=-\mathbb{E}(Y,(Z-b))$ implying that $\mathbb{E}(Y(Z-b))=0$ from which the result follows. ∎
Proposition 2.
If a random vector $(Y,Z)^{\top}$ has continuous marginal distributions and is either h-symmetric or v-symmetric, then the Spearman rank correlation $\rho_{S}(Y,Z)$ and the Kendall rank correlation $\tau(Y,Z)$ are both zero.
Proof.
Again it suffices to give the proof in the case of h-symmetry. Let the unique copula of $(Y,Z)^{\top}$ be denoted $C$. For Spearman’s rank correlation we need to calculate $\rho_{S}(Y,Z)=12\int_{0}^{1}\int_{0}^{1}(C(u,v)-uv)\mathrm{d}u\mathrm{d}v$. From the identity $C(u,v)=u-C(u,1-v)$ we infer that
$\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v$ $=$ $\int_{0}^{1}\int_{0}^{1}u\mathrm{d}u\mathrm{d}v-\int_{0}^{1}\int_{0}^{1}C(u,1-v)\mathrm{d}u\mathrm{d}v=\frac{1}{2}-\int_{0}^{1}\int_{0}^{1}C(u,y)\mathrm{d}u\mathrm{d}y$
where we have made the change of variable $y=1-v$. We must have $\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v=\tfrac{1}{4}$ and then simple calculations give $\rho_{S}(Y,Z)=0$.
In the case of Kendall’s rank correlation we need to compute $\tau(Y,Z)=4\mathbb{E}(C(U,V))-1$ where $(U,V)^{\top}$ is a pair of variables with distribution function $C$. The identity $C(u,v)=u-C(u,1-v)$ implies that $(U,V)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,1-V)^{\top}$; indeed $(U,V-0.5)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,0.5-V)^{\top}$ since $(U,V)^{\top}$ is h-symmetric about 0.5. Hence we infer that
$\mathbb{E}(C(U,V))$ $=$ $\mathbb{E}(U)-\mathbb{E}(C(U,1-V))=\frac{1}{2}-\mathbb{E}(C(U,V))$
which implies that $\mathbb{E}(C(U,V))=\tfrac{1}{4}$ and $\tau(Y,Z)=0$. ∎
Despite the absence of correlation in distributions with these forms of symmetry, we will see in this paper that they can have quite strong forms of dependence.
2.3 Copula densities
We will analyse GARCH-type models in terms of their joint densities. Henceforth we assume that the generic random vector $(Y,Z)^{\top}$ has joint density $f_{Y,Z}$. It is easily inferred from Definition 1, part (i), that h-symmetry is equivalent to $f_{Y,Z}(y,b+z)=f_{Y,Z}(y,b-z)$ for all $(y,z)\in\mathbb{R}^{2}$ and some $b\in\mathbb{R}$. Similarly, v-symmetry and radial symmetry are equivalent to the density satisfying $f_{Y,Z}(a+y,z)=f_{Y,Z}(a-y,z)$ for some $a\in\mathbb{R}$ and $f_{Y,Z}(a+y,b+z)=f_{Y,Z}(a-y,b-z)$ for some $a,b\in\mathbb{R}$, respectively.
The density $c$ of the copula $C$ of $(Y,Z)^{\top}$ is given by
$c(u,v)=\frac{f_{Y,Z}\left(F_{Y}^{-1}(u),F_{Z}^{-1}(v)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{Z}\left(F_{Z}^{-1}(v)\right)}$ (3)
where $f_{Y}$ and $f_{Z}$ denote the marginal densities. The identity $C(u,v)=u-C(u,1-v)$ for h-symmetry of the copula $C$ in Definition 2 is equivalent to the identity $c(u,v)=c(u,1-v)$ for densities (symmetry in the line $v=0.5$). Similarly, v-symmetry is equivalent to $c(u,v)=c(1-u,v)$ (symmetry in the line $u=0.5$) and radial symmetry is equivalent to $c(u,v)=c(1-u,1-v)$.
The following result provides the first connection between copula densities and v-transforms. The function $\mathcal{V}(u)=|2u-1|$ appearing in this result is the first and simplest example of a v-transform, as we will discuss in Section 4.
Proposition 3.
Let $(Y,Z)^{\top}$ be a random vector with joint density $f_{Y,Z}$ and copula density $c$. Then, for all $(u,v)\in[0,1]^{2}$,
  1. (i)
    if $(Y,Z)^{\top}$ is h-symmetric about $b$, $c(u,v)=c^{*}(u,|2v-1|)$ where $c^{*}$ is the density of the copula of $(Y,|Z-b|)^{\top}$;
  2. (ii)
    if $(Y,Z)^{\top}$ is v-symmetric about $a$, $c(u,v)=c^{*}(|2u-1|,v)$ where $c^{*}$ is the density of the copula of $(|Y-a|,Z)^{\top}$.
Proof.
It suffices to show part (i). We set $\tilde{Z}=Z-b$ and observe that $(Y,\tilde{Z})^{\top}$ is h-symmetric about 0. We then find that
$F_{Y,|\tilde{Z}|}(y,z)=\mathbb{P}(Y\leqslant y,|\tilde{Z}|\leqslant z)=\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant z)-\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant-z)=F_{Y,\tilde{Z}}(y,z)-F_{Y,\tilde{Z}}(y,-z)$
and hence $f_{Y,|\tilde{Z}|}(y,z)=f_{Y,\tilde{Z}}(y,z)+f_{Y,\tilde{Z}}(y,-z)=2f_{Y,\tilde{Z}}(y,z)$ by h-symmetry about zero. Similarly, $f_{|\tilde{Z}|}(z)=2f_{\tilde{Z}}(z)$. From (3) we have
$c^{*}(u,|2v-1|)=\frac{f_{Y,|\tilde{Z}|}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{|\tilde{Z}|}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}=\frac{f_{Y,\tilde{Z}}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{\tilde{Z}}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}.$
Since $F_{|\tilde{Z}|}(z)=2F_{\tilde{Z}}(z)-1$ for $z\geqslant 0$, it follows that $F_{|\tilde{Z}|}^{-1}(q)=F_{\tilde{Z}}^{-1}((1+q)/2)$ for $q\in[0,1]$. Therefore, for $v\leqslant 0.5$, we have
$F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(1-2v)=F_{\tilde{Z}}^{-1}(1-v)=-F_{\tilde{Z}}^{-1}(v)$
while, for $v\geqslant 0.5$, we have $F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(2v-1)=F_{\tilde{Z}}^{-1}(v)$. Using the h-symmetry of $f_{Y,\tilde{Z}}$ and the symmetry of $f_{\tilde{Z}}$ we infer that $c^{*}(u,|2v-1|)=\tilde{c}(u,v)$ where $\tilde{c}$ is the density of the copula of $(Y,\tilde{Z})^{\top}$. The identity $c^{*}(u,|2v-1|)=c(u,v)$ follows from the fact that $(Y,Z)^{\top}$ and $(Y,\tilde{Z})^{\top}$ have the same copula. ∎
Clearly, by putting both parts of Proposition 3 together, we see that the density of the copula of a random vector $(Y,Z)^{\top}$ that is jointly symmetric about $(a,b)$ satisfies $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|Y-a|,|Z-b|)^{\top}$.
3 ARCH and GARCH copulas
3.1 ARCH and GARCH processes
We define a broad class of first-order GARCH-type models and establish nomenclature for some special cases.
Definition 3 (First-order GARCH processes).
  1. 1.
    A process of GARCH(1,1) type is a strictly stationary process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations of the form
    $X_{t}=\sigma_{t}\;\epsilon_{t}$ (4)
    where $(\epsilon_{t})_{t\in\mathbb{Z}}$ is a sequence of iid innovations, drawn from a continuous distribution with mean zero and variance one, and $(\sigma_{t})_{t\in\mathbb{Z}}$ is a sequence of variables satisfying $\sigma_{t}=\sigma(X_{t-1},\sigma_{t-1})$ where $\sigma:\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a strictly positive-valued parametric volatility function.
  2. 2.
    A process of ARCH(1) type is a process of GARCH(1,1) type where the volatility function $\sigma(x,s)$ depends only on $x$.
  3. 3.
    The classical GARCH(1,1) process is a process of GARCH(1,1) type type in which the volatility function satisfies
    $\sigma(x,s)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}+\beta_{1}s^{2}}$ (5)
    where $\alpha_{0}>0$, $\alpha_{1}\geqslant 0$, $\beta_{1}\geqslant 0$ and $\mathbb{E}(\ln(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}))<0$ for an innovation $\epsilon_{1}$.
  4. 4.
    The classical ARCH(1) process is the classical GARCH(1,1) process with $\beta_{1}=0$.
  5. 5.
    A process of GARCH(1,1) type is said to have symmetric distribution, if the distribution of $\epsilon_{1}$ has a density that is symmetric around zero, and to have symmetric volatility if $\sigma(x,w)$ is an even function of $x$. It is said to be symmetric if it has symmetric distribution and symmetric volatility.
The mathematical properties of the classical GARCH(1,1) process have been widely studied and the conditions on the parameters in part 3 of Definition 3 are known to guarantee a strictly stationary solution to the equations (4(Nelson,, 1990; Bougerol and Picard,, 1992). For a finite-variance solution it is usual to assume the stronger condition that $\alpha_{1}+\beta_{1}<1$. It is known that the marginal or stationary distribution $F_{X}$ has regularly varying (power-law) tails and Mikosch and Stărică, (2000) showed that the tail index $\zeta$ may be calculated as the root of the equation $\mathbb{E}\left(\left(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}\right)^{\zeta/2}\right)=1$. However, the exact functional forms of $F_{X}$ or its density $f_{X}$ are not available.
3.2 The first-order copula of GARCH(1,1)-type processes
Simple calculations show that the joint density of a pair of successive variables $(X_{1},X_{2})^{\top}$ in a process of GARCH(1,1) type is given by
$f_{X_{1},X_{2}}(x,y)=\int_{0}^{\infty}f_{X_{1},X_{2},\sigma_{1}}(x,y,s)\mathrm{d}s$ $=\int_{0}^{\infty}f_{X_{2}\mid X_{1},\sigma_{1}}(y\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
$=\int_{0}^{\infty}\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{y}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s$ (6)
where $f_{\sigma}$ is the marginal density of the volatility process $(\sigma_{t})_{t\in\mathbb{Z}}$ and $f_{\epsilon}$ is the density of the innovations. Applying (3), the density $c_{1}$ of the copula $C_{1}$ of $(X_{1},X_{2})^{\top}$ is thus
$c_{1}(u,v)=\frac{\int_{0}^{\infty}\frac{1}{s\sigma\left(F_{X}^{-1}(u),s\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)}{\sigma\left(F_{X}^{-1}(u),s\right)}\right)f_{\epsilon}\left(\frac{F_{X}^{-1}(u)}{s}\right)f_{\sigma}(s)\mathrm{d}s}{f_{X}\left(F_{X}^{-1}(u)\right)f_{X}\left(F_{X}^{-1}(v)\right)}$ (7)
where
$f_{X}(x)=\int_{0}^{\infty}f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s=\int_{0}^{\infty}\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s$ (8)
is the marginal density of the process $(X_{t})_{t\in\mathbb{Z}}$.
In the case of a process of ARCH(1) type (where $\sigma(x,s)$ only depends on $x$) these expressions simplify considerably and become
$f_{X_{1},X_{2}}(x,y)$ $=\frac{1}{\sigma(x)}f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)f_{X}(x),\quad\quad c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$ (9)
The key insight from these calculations is the following.
Proposition 4.
  1. (i)
    In a GARCH(1,1)-type process with symmetric distribution, the distribution of $(X_{1},X_{2})^{\top}$ is h-symmetric about zero.
  2. (ii)
    In a symmetric GARCH(1,1)-type process, the distribution of $(X_{1},X_{2})^{\top}$ is jointly symmetric about the origin.
Proof.
We consider the joint density $f_{X_{1},X_{2}}(x,y)$ in (6).
  1. (i)
    It is clear that a symmetric innovation distribution implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(x,-y)$, meaning the random vector $(X_{1},X_{2})^{\top}$ is h-symmetric about zero.
  2. (ii)
    An even volatility function also implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(-x,y)$ and so $(X_{1},X_{2})^{\top}$ is v-symmetric about zero and hence radially and jointly symmetric about the origin.
In case (i) above we can apply Lemma 1 to infer that the copula $C_{1}$ is h-symmetric and we can apply Proposition 3 to infer that its density may be written in the form $c_{1}(u,v)=c_{X_{1},|X_{2}|}(u,|2v-1|)$ where $c_{X_{1},|X_{2}|}$ is the density of the copula of $(X_{1},|X_{2}|)^{\top}$. In case (ii) we infer that the copula $C_{1}$ is jointly symmetric and its density may be written in the form $c_{1}(u,v)=c_{|X_{1}|,|X_{2}|}(|2u-1|,|2v-1|)$ where $c_{|X_{1}|,|X_{2}|}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
Despite the strong symmetry revealed by Proposition 4, there is one important form of symmetry that symmetric processes of GARCH(1,1) type do not have in general, namely exchangeability. It is not true that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(y,x)$ for all $(x,y)^{\top}\in\mathbb{R}^{2}$ so the bivariate distributions of the pairs $(X_{1},X_{2})^{\top}$ and $(X_{2},X_{1})^{\top}$ differ. This implies that ARCH and GARCH are not time-reversible models. In general $f_{X_{2}\mid X_{1}}(y\mid x)\neq f_{X_{1}\mid X_{2}}(y\mid x)$ and thus, in an intuitive sense, they are models in which the predictive distribution of the future given the past differs from the “predictive distribution” of the past given the future. From an econometric point of view, this would be considered natural. The non-exchangeability carries over to the copula so that we have in general that $C_{1}(u,v)\neq C_{1}(v,u)$ on $[0,1]^{2}$. Despite this observation, it is notable that the majority of the copulas employed in vine copula models of financial time series are exchangeable ones.
3.3 Computational examples
In this section we use computational methods to approximate and plot the joint densities $f_{X_{1},X_{2}}$ and copula densities $c_{1}$ in a number of models of GARCH(1,1) type. We concentrate on processes of ARCH(1) type in our illustrations since their first-order Markov structure simplifies computation of the unknown marginal density $f_{X}$ and distribution function $F_{X}$, which are key ingredients in (9). Beginning with the equation
$f_{X}(x)=\int_{-\infty}^{\infty}f_{\epsilon}\left(\frac{x}{\sigma(y)}\right)\frac{1}{\sigma(y)}f_{X}(y)\mathrm{d}y,$ (10)
which follows easily from (9), we approximate the integral by a sum of weighted values of $f_{X}(y_{j})$ on a grid of points $\mathcal{Y}=\{y_{1},\ldots,y_{n}\}$; we can then solve a linear system for $f_{X}(y_{j}),y_{j}\in\mathcal{Y}$, which is accurate enough for our purposes.
We also consider the classical GARCH(1,1) model with Gaussian innovations. To analyse this process we use simulation and estimation to find a parametric model for the unknown marginal density $f_{\sigma}$ of the volatility process. We can then integrate out volatility in expressions like (6), (7) and (8) [^1] All details of the model for $f_{\sigma}$ and the discretization method for ARCH(1)-type processes can be found in the code that accompanies this paper..
The parameter values for all the models we consider are summarized in Table 1.
Process $\alpha_{0}$ $\alpha_{1}$ $\beta_{1}$ $\gamma_{1}$
ARCH(1) 0.4 0.6 0 0
ARCH(1) with leverage effect 0.4 0.3 0 0.4
GARCH(1,1) 0.1 0.3 0.6 0
GARCH(1,1) with leverage effect 0.1 0.1 0.6 0.4
Table 1: Parameter values for processes analysed in paper. The volatility function $\sigma(x,s)$ in all models satisfies $\sigma(x,s)^{2}=\alpha_{1}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}+\beta_{1}s^{2}$. Innovation distributions used are either Gaussian, Student t with $\nu=4$ or $\nu=2.5$ degrees of freedom or skewed Student t with $\nu=4$ degrees of freedom and skewness parameter $\lambda=0.8$; all innovation distributions are scaled to have mean zero and variance one.
Figure 1: Contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values. Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations. Bottom right: classical GARCH(1,1) process with Gaussian innovations.
Figure 2: Contour plots of $c_{1}(u,v)$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values. Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations. Bottom right: classical GARCH(1,1) process with Gaussian innovations.
  1. 1.
    Classical ARCH(1) process with Gaussian innovations (symmetric) and parameter values $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$ as in Table 1; this process has unit variance and tail index $\zeta=3.82$. The joint density $f_{X_{1},X_{2}}$ and copula density $c_{1}$ are shown in the top left panels of Figures 1 and 2 respectively. The non-exchangeability, evidenced by a lack of symmetry in the dashed lines $x_{2}=x_{1}$ and $v=u$, is evident in both pictures. We can also observe the joint symmetry, that is the rotational symmetry of order 2 as well as the h-symmetry and v-symmetry.
    Changing $\alpha_{1}$ changes the degree of dependence in the pictures. In particular, as $\alpha_{1}\to 0$ the copula converges to the independence copula while increasing $\alpha_{1}$ serves to strengthen the cross-shaped character of the copula density. For normal innovations, the theoretical maximum value of $\alpha_{1}$ to obtain a strictly stationary process is approximately 3.56, but for values $\alpha_{1}\geqslant 1$ we obtain infinite-variance processes which are seldom used in practical applications and become increasingly pathological with increasing $\alpha_{1}$ as their tail indexes converge to zero (Mikosch and Stărică,, 2000).
  2. 2.
    Classical ARCH(1) process with Student innovations (symmetric). This is shown in the top right panels of Figures 1 and 2 for a process with the same parameters as in the previous example and degree of freedom $\nu=4$ for the innovation distribution; the tail index of the process is 2.76. Remarkably, the resulting joint density appears to be very close to spherical and the copula appears to be very close to an exchangeable copula for this combination of parameters; the joint symmetry is again apparent. The bottom left panels of Figures 1 and 2 show a process in which $\nu=2.5$ and the tail index is 2.09; these show non-exchangeability more clearly.
    It is notable that, although the tail index of the marginal distribution increases with decreasing degree of freedom (increasing tail index of the innovation density), the dependence actually weakens. We can estimate a distance $D_{1}=\int_{0}^{1}\int_{0}^{1}|c_{1}(u,v)-1|\mathrm{d}u\mathrm{d}v$ between the copula density $c_{1}$ and the density of the independence copula; for the Gaussian, t4 and t2.5 copulas we obtain respectively 0.213, 0.158 and 0.092.
  3. 3.
    Classical GARCH(1,1) process with Gaussian innovations (symmetric). This is shown in the bottom right panels of Figures 1 and 2 for a process with Gaussian innovations and parameters as in Table 1, which has variance one and tail index 4.09.
    It is well known that, for a covariance stationary GARCH(1,1) process ($\alpha_{1}+\beta_{1}<1$), the squared process $(X_{t}^{2})_{t\in\mathbb{Z}}$ behaves like an ARMA(1,1) process with autoregressive parameter $\alpha_{1}+\beta_{1}$ and moving-average parameter $-\beta_{1}$; see, for example, McNeil et al., (2015, Section 4.2.2). Thus the value of $\alpha_{1}+\beta_{1}$ controls the persistence, or the rate at which serial dependencies in the squared (or absolute value) process die away with lag. For a fixed value of $\alpha_{1}+\beta_{1}$ the magnitude of the serial dependency at lag 1 increases with $\alpha_{1}$, so that the ARCH(1) special case actually has the strongest dependence and the copula tends towards independence as $\alpha_{1}\to 0$. The parameters we have chosen are not necessarily typical of real financial data, where the ARCH effect $\alpha_{1}$ might be smaller and the GARCH effect $\beta_{1}$ larger, but have instead been selected to give a picture that has more evident dependence. The main point is to show that the behaviour of the GARCH(1,1) copula is qualitatively similar to that of the ARCH(1) copula.
  4. 4.
    ARCH(1) model with Gaussian innovations and leverage (symmetric distribution). Here we use a volatility function $\sigma(x)$ satisfying $\sigma(x)^{2}=\alpha_{0}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}$ for $\gamma_{1}>0$ so that a negative return produces a larger value of the volatility than a positive return of the same magnitude. From now on, we concentrate on contour plots of the copula only, as shown in the top left panel of Figure 3. It can be seen that leverage removes radial symmetry and v-symmetry, but there is still h-symmetry (symmetry in the line $v=0.5$).
  5. 5.
    Classical ARCH(1) model with skewed Student innovations (asymmetric distribution but symmetric volatility). This is shown in the top right panel of Figure 3 for an innovation distribution that is skewed to the left (towards negative shocks). The skewness removes all the previous symmetries in the pictures.
  6. 6.
    Absolute value copula for classical ARCH(1) model with Gaussian innovations. This is shown in the bottom left of Figure 3 for the same process as in the top left of Figure 2. It can be thought of as a blown-up version of the top right quadrant of the latter and clearly shows non-exchangeability and stronger dependence in the upper tail. Note that this picture also illustrates the density of the volatility copula of ARCH(1), that is the copula of $(\sigma_{1},\sigma_{2})^{\top}$; this follows because $\sigma_{t}$ is a strictly increasing function of $|X_{t-1}|$ for all $t$ and copulas are invariant under strictly increasing transformations of variables. For the same reason, it is also the copula density for the variables $(X_{1}^{2},X_{2}^{2})^{\top}$.
Figure 3: Introducing asymmetry. Top left: ARCH(1) with Gaussian innovations in which the volatility function has leverage and the parameters are $\alpha_{0}=0.4$, $\alpha_{1}=0.3$ and $\gamma_{1}=0.4$. Top right: classical ARCH(1) model with with 4 degrees of freedom and skewness parameter $\lambda=0.8$ where $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. Bottom left: copula density for copula of the absolute values for classical ARCH(1) copula with Gaussian innovations. Bottom right: AR(1)-ARCH(1) model with Gaussian innovations and parameters $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
3.4 AR(1) model with ARCH(1) errors
We consider briefly a time series model with stochastic volatility as well as a non-zero stochastic conditional mean term. We consider an AR(1) model with ARCH(1) errors and show the implications of the conditional mean term. Specifically, we consider the first-order Markov process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations
$X_{t}=\phi X_{t-1}+\sigma(X_{t-1})\;\epsilon_{t}\quad\text{with}\quad\sigma(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}},$ (11)
where the innovations $(\epsilon_{t})_{t\in\mathbb{Z}}$ have a symmetric distribution. This model was proposed and analysed by Borkovec and Klüppelberg, (2001) [^2] It differs from the usual version of the AR(1)-ARCH(1) model, $X_{t}=\phi X_{t-1}+\sigma(X_{t-1}-\phi X_{t-2})\;\epsilon_{t}$, but this is not first-order Markov and thus less straightforward to treat with our methods. who showed that it is covariance stationary provided $\phi^{2}+\alpha_{1}<1$ and has a symmetric stationary distribution $F_{X}$.
The formulas (9) require only simple modification to account for the non-zero conditional mean term. In particular, the copula becomes
$c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)-\phi F_{X}^{-1}(u)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$
It may be easily verified that the copula density satisfies $c(u,v)=c(1-u,1-v)$ and so the copula is radially symmetric. However, the copula is neither h-symmetric nor v-symmetric, and thus it is not jointly symmetric. The copula density is shown in the bottom right panel of Figure 3 in the case where $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. It is evident that more of the probability mass is now on the main diagonal of the copula.
In the remainder of the paper we concentrate on models without a conditional mean term, but this example gives some clues about the effect on the copula of serial dependence in the untransformed process $(X_{t})_{t\in\mathbb{Z}}$ in combination with stochastic volatility, a point we return to in our discussion in Section 7.
3.5 Higher-order conditional copulas of GARCH(1,1)-type processes
For higher-order copulas we introduce the compact notation $\bm{X}_{[j,l]}=(X_{j},\ldots,X_{l})^{\top}$ to denote random vectors consisting of a contiguous block of random variables; we apply the same notation to realisations of the random vectors. Whenever the notation is used with $j>l$, this means that the variables are taken in reverse order to their natural temporal order.
Now we consider the copula $C_{k}$ of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $\bm{X}_{[2:k]}=\bm{x}_{[2:k]}$ for $k\geqslant 2$. The density of the conditional copula takes the form
$c_{k}(u,v\mid\bm{x}_{[2:k]})=\frac{f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right),F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}{f_{X_{1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)f_{X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}$
Note that this copula density will simply be the independence copula density for a process of ARCH(1) type, since $X_{1}$ and $X_{3}$ are conditionally independent given $X_{2}$. Consider, therefore a process with an explicit GARCH effect, i.e. $\sigma(x,s)$ depends also on $s$.
To develop intuition we consider the case where $k=2$. The symmetry properties can be investigated by considering the joint conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$. To this end consider the joint density
$f_{X_{1},X_{2},X_{3}}(x,z,y)$ $=\int_{0}^{\infty}f_{X_{1},X_{2},X_{3},\sigma_{1}}(x,z,y,s)\mathrm{d}s$
$=\int_{0}^{\infty}f_{X_{3}\mid X_{2},X_{1},\sigma_{1}}(y\mid z,x,s)f_{X_{2}\mid X_{1},\sigma_{1}}(z\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
$=\int_{0}^{\infty}\frac{1}{\sigma(z,\sigma(x,s))}f_{\epsilon}\left(\frac{y}{\sigma(z,\sigma(x,s))}\right)\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{z}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s.$
It is straightforward to see that a process with symmetric innovations will satisfy $f_{X_{1},X_{2},X_{3}}(x,z,-y)=f_{X_{1},X_{3},X_{2}}(x,z,y)$ while a symmetric process will also satisfy $f_{X_{1},X_{2},X_{3}}(-x,z,y)=f_{X_{1},X_{2},X_{3}}(x,z,y)$. These properties will also hold for the conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)=f_{X_{1},X_{2},X_{3}}(x,z,y)/f_{X}(z)$ and thus, by Lemma 1, the various forms of symmetry will be inherited by the conditional copula density $c_{2}(u,v\mid z)$. By applying a similar analysis to the joint density $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)$ we can therefore establish the following result.
Proposition 5.
  1. (i)
    In a GARCH(1,1)-type process with symmetric distribution, the conditional copula $C_{k}$ is h-symmetric about zero for $k\geqslant 2$.
  2. (ii)
    In a symmetric GARCH(1,1)-type process the conditional copula is jointly symmetric about the origin for $k\geqslant 2$.
Proof.
Suppose we write $s_{2}(x,s)=\sigma(x,s)$, $s_{3}(z_{1},x,s)=\sigma(z_{1},s_{2}(x,s))$ and then for $j=4,\ldots,k+1$ we define recursive functions $s_{j}(z_{j-2},\ldots,z_{1},x,s)=\sigma(z_{j-2},s_{j-1}(z_{j-3},\ldots,z_{1},x,s))$. This permits us to write the joint density as
$f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)=\int_{0}^{\infty}f_{X_{k+1}\mid\sigma_{k+1}}(y\mid s_{k+1}(z_{k-1},\ldots,z_{1},x,s))\\ \left(\prod_{j=3}^{k}f_{X_{j}\mid\sigma_{j}}(z_{j-1}\mid s_{j}(z_{j-2},\ldots,z_{1},x,s))\right)f_{X_{2}\mid\sigma_{2}}(z_{1}\mid s_{2}(x,s))f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
It is clear that this joint density satisfies $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(x,\ldots,-y)$ when the innovation distribution is symmetric. Moreover it is also clear that in the symmetric model we have $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(-x,\ldots,y)$ since all of the functions $s_{2},\ldots,s_{k+1}$ are even functions of $x$. These properties will also hold for the conditional density
$f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}(x,y\mid z_{1},\ldots,z_{k-1})=f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)/f_{X_{2},\ldots,X_{k}}(z_{1},\ldots,z_{k-1})$
and thus we can apply Lemma 1 to infer that the assertions regarding symmetry of the copula hold. ∎
Remark 1.
Since the joint density of $(X_{1},X_{k+1})^{\top}$ can be obtained as
$f_{X_{1},X_{k+1}}(x,y)=\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)\mathrm{d}z_{1}\cdots\mathrm{d}z_{k-1}$
it follows that the same method of proof can be used to show that the copula of the unconditional distribution of $(X_{1},X_{k+1})^{\top}$ shares the same symmetry properties as the conditional copula.
An interesting question concerns the degree to which $c_{k}(u,v\mid\cdot)$ depends on the realized values of the conditioning variables. We investigate this question for $k=2$ and and write $c_{2}(u,v;w):=c_{2}(u,v\mid F_{X}^{-1}(w))$. The latter function takes the form
$c_{2}(u,v;w)=\frac{f_{X_{1},X_{3}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right),F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}{f_{X_{1}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)f_{X_{3}\mid X_{2}}\left(F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}$
and can be calculated approximately for the classical GARCH(1,1) model building on our analysis in Section 3.3.
In Figure 4 we show contour plots for the cases where $w=0.5$, i.e. the intermediate observation is at the median of the marginal distribution (zero in this case), and $w=0.75$, i.e. the intermediate observation is at a quartile of the marginal distribution. While the two pictures are not identical, they are relatively similar to each other in terms of the shapes and heights of the contour lines. To examine this further we attempt to compute an approximate value for the absolute distance between the conditional copula densities and the independence copula density using the measure $D_{2}(w)=\int_{0}^{1}\int_{0}^{1}|c_{2}(u,v;w)-1|\mathrm{d}u\mathrm{d}v$. We obtain the estimates $D_{2}(0.5)=0.10$ and $D_{2}(0.75)=0.09$, which can be compared with the equivalent measure $D_{1}=0.16$ for the copula density $c_{1}$ in Figure 2.
As might be expected, the dependence weakens from lag 1 to lag 2. Moreover, the value of the conditioning variable does not seem to have a particularly large effect on the conditional copula at lag 2. This might be taken as weak evidence that the simplifying assumption is approximately tenable; obviously we should look at values of $w$ further into the tail, but our approximation methods then become more unreliable.
Figure 4: Contour plots of $c_{2}(u,v;w)$ for the classical GARCH(1,1) process with Gaussian innovations. In the left-hand figure $w=0.5$ and in the right-hand figure $w=0.75$.
4 Copula densities and v-transforms
4.1 V-transforms and their inverses
The function $\mathcal{V}(u)=|2u-1|$ which appeared in Proposition 3 is a symmetric v-shaped function and the simplest example of a v-transform. These are special cases of uniform-measure-preserving transformations (Porubský et al.,, 1988) and are extensively analysed in McNeil, (2021) and Bladt and McNeil, 2022b . In this section we recap the essential elements of the theory and examine its relevance for understanding and mimicking the behaviour of the copula of processes of GARCH(1,1) type.
Definition 4.
A mapping $\mathcal{V}:[0,1]\to[0,1]$ is a v-transform if it takes the form
$\mathcal{V}(u)=\begin{cases}(1-u)-(1-\delta)\Psi\left(\frac{u}{\delta}\right)&u\leqslant\delta,\\ u-\delta\Psi^{-1}\left(\frac{1-u}{1-\delta}\right)&u>\delta,\end{cases}$ (12)
for a parameter $10<\delta<$ 1, known as the fulcrum, and a continuous and strictly increasing distribution function $\Psi$ on $[0,1]$, known as the generator.
Mappings of this form are always v-shaped with $\mathcal{V}(0)=\mathcal{V}(1)=1$ and $\mathcal{V}(\delta)=0$. They preserve uniformity in the sense that if $U\sim U(0,1)$, then $\mathcal{V}(U)\sim U(0,1)$. If the generator $\Psi$ is taken to be standard uniform, we obtain the linear v-transform, which satisfies $\mathcal{V}(u;\delta)=(\delta-u)/\delta$ for $u\leqslant\delta$ and $\mathcal{V}(u;\delta)=(u-\delta)/(1-\delta)$ for $u>\delta$; this includes the special case $\mathcal{V}(u;0.5)=|2u-1|$, which is the symmetric v-transform. It is possible to introduce further parameters. For example, by taking $\Psi(x)=x^{\kappa}$ for $\kappa>0$, we obtain a two-parameter family
$\mathcal{V}(u;\delta,\kappa)=\begin{cases}(1-u)-(1-\delta)\left(\frac{u}{\delta}\right)^{\kappa}&u\leqslant\delta,\\ u-\delta\left(\frac{1-u}{1-\delta}\right)^{1/\kappa}&u>\delta.\end{cases}$ (13)
For further results we need some more notation. For every $y\in(0,1]$ there are two possible values $u$ on either side of the fulcrum $\delta$ such that $\mathcal{V}(u)=y$ and these may be written as $\mathcal{V}^{-1}(y)$ and $\mathcal{V}^{-1}(y)+y$ where $\mathcal{V}^{-1}(y)=\inf\{u:\mathcal{V}(u)=y\}$ is a partial inverse of $\mathcal{V}$. We also introduce the function $\Delta(y)=-1/\mathcal{V}^{\prime}(\mathcal{V}^{-1}(y))$ where $\mathcal{V}^{\prime}$ is the derivative of $\mathcal{V}$. If $U\sim U(0,1)$ then McNeil, (2021) shows that $\Delta(y)=\mathbb{P}(U\leqslant\delta\mid\mathcal{V}(U)=y)$ for $y\in(0,1]$; for the linear v-transform $\Delta(y)=\delta$, but otherwise $\Delta(y)$ depends explicitly on $y$.
We can stochastically invert a v-transform in such a way that we also preserve uniformity of random variables. For $y\in(0,1]$, the stochastic inverse $\mathcal{V}^{\leftarrow}(y)$ of the v-transform $\mathcal{V}$ is defined to take the value $\mathcal{V}^{-1}(y)$ (from the left branch of the transform) with probability $\Delta(y)$ and the value $\mathcal{V}^{-1}(y)+y$ (from the right branch of the transform) with probability $1-\Delta(y)$; for $y=0$ we set $\mathcal{V}^{\leftarrow}(y)=\delta$. If $U\sim U(0,1)$ then it can also be shown that $\mathcal{V}^{\leftarrow}(U)\sim U(0,1)$. Note, however, that while $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ holds for all $y\in[0,1]$, the equation $\mathcal{V}^{\leftarrow}\circ\mathcal{V}(u)=u$ is only true with probability $\Delta(u)$, if $u<\delta$, and with probability $1-\Delta(u)$, if $u>\delta$.
4.2 V-transformation and inverse v-transformation of copulas
The following result, which is an adaptation of a result in McNeil, (2021), is key to constructing new copulas using v-transforms and their inverses.
Theorem 1.
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be v-transforms with associated partial inverses $\mathcal{V}_{1}^{-1}$ and $\mathcal{V}_{2}^{-1}$ and $\Delta$-functions $\Delta_{1}$ and $\Delta_{2}$. Let $(U,V)^{\top}$ be distributed according to a copula $C_{U,V}$ with density $c_{U,V}$.
  1. (i)
    $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ with density
    $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)=\;$ $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v))\;\Delta_{1}(u)\,\Delta_{2}(v)\;+$
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v))\;\big{(}1-\Delta_{1}(u)\big{)}\,\Delta_{2}(v)\;+$ (14)
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v)+v)\;\Delta_{1}(u)\,\big{(}1-\Delta_{2}(v)\big{)}\;+$
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v)+v)\;\big{(}1-\Delta_{1}(u)\big{)}\,\big{(}1-\Delta_{2}(v)\big{)}.$
  2. (ii)
    $(\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}^{\leftarrow}_{1}(U),\mathcal{V}^{\leftarrow}_{2}(V)}$ with density
    $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}(u,v)=c_{U,V}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v)).$ (15)
Proof.
This may be proved by a simple adaptation of the proof of Theorem 3 in McNeil, (2021) which deals with the case of two (or more) identical v-transforms. ∎
We refer to the copula density $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ in part (i) as a v-transformed version of $c_{U,V}$ while we refer to the copula density $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}$ in part (ii) as an inverse-v-transformed version of $c_{U,V}$. Inverse v-transformation of a copula density using (15) followed by v-transformation using (14) returns the original copula density $c_{U,V}$ (since for any v-transform $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ for $y\in[0,1]$). However v-transformation of a copula density using (14) followed by inverse v-transformation using (15) does not, in general, return the original density $c$, except in special cases, such as the following.
Corollary 1.
If $C_{U,V}$ is a jointly symmetric copula and $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$ for $u\in[0,1]$, then the density of the copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ satisfies
$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=$ $c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)$ (16)
$=$ $c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)$
and
$c_{\mathcal{V}_{1}^{\leftarrow}\circ\mathcal{V}_{1}(U),\mathcal{V}_{2}^{\leftarrow}\circ\mathcal{V}_{2}(V)}(u,v)=c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(|2u-1|,|2v-1|)=c_{U,V}(u,v).$ (17)
Proof.
By part (i) of Theorem 1 and equation (14), the copula density satisfies
$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=\frac{1}{4}\left(c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)\right.$
$\left.+c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)\right)$
and (16) follows easily from the joint symmetry property of $C_{U,V}$. The identity (17) is then easily verified by considering the values of $|2u-1|$ and $|2v-1|$ in the four quadrants of the unit square defined by the lines $u=0.5$ and $v=0.5$ and using each of the four equations in (16) in turn. ∎
Remark 2.
We can easily derive a corresponding result for the simpler case where the copula $C_{U,V}$ has h-symmetry and $V$ is v-transformed with the symmetric v-transform. The copula density of the random vector $(U,|2V-1|))^{\top}$ can be shown to satisfy
$c_{U,|2V-1|}(u,v)=c_{U,V}\left(u,\frac{1-v}{2}\right)=c_{U,V}\left(u,\frac{1+v}{2}\right)$ (18)
and it is easily verified that $c_{U,|2V-1|}(u,|2v-1|)=c_{U,V}(u,v)$.
Remark 3.
To see that the identity (17) does not always hold when $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$, it suffices to start with any copula $C_{U,V}$ that is not jointly symmetric, since application of the inverse v-transform in the final step always results in a jointly symmetric copula. For example, when $C_{U,V}=C^{\text{Ga}}_{\rho}$, the Gaussian copula with parameter $\rho>0$, v-transformation followed by inverse v-transformation results in the jointly symmetric mixture $0.5C^{\text{Ga}}_{\rho}+0.5C^{\text{Ga}}_{-\rho}$.
Corollary 1 and Remark 2 allow us to state the following simple result, which is relevant to our study of the symmetries in the densities of copulas of GARCH(1,1) type.
Proposition 6.
Let $C$ be an absolutely continuous copula with density $c$.
  1. (i)
    $C$ is h-symmetric if and only if $c(u,v)=c^{*}(u,|2v-1|)$ for some copula density $c^{*}$.
  2. (ii)
    $C$ is jointly symmetric if and only if $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some copula density $c^{*}$.
Proof.
It is trivial to verify that copulas with the structures $c^{*}(u,|2v-1|)$ and $c^{*}(|2u-1|,|2v-1|)$ are h-symmetric and jointly symmetric, respectively. To show the reverse implication in part (i), we choose $c^{*}$ to be the copula density of $(U,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$. To show the reverse implication in part (ii), we choose $c^{*}$ to be the copula density of $(|2U-1|,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$. ∎
4.3 Implicit v-transforms in general bivariate densities
In this section we use part (i) of Theorem 1 to find a condition for our generic random vector $(Y,Z)^{\top}$ on $\mathbb{R}^{2}$ which allows us to write its copula density $c$ in the general form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ for v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ and a copula density $c^{*}$. This result requires the concept of a v-shaped transformation on $\mathbb{R}$, which should be distinguished from a v-transform.
Definition 5.
A function $T$ is a v-shaped function on $\mathbb{R}$ if it can be written in the form
$T(x)=\begin{cases}T_{L}(\mu_{T}-x)&\quad x\leqslant\mu_{T}\\ T_{R}(x-\mu_{T})&\quad x>\mu_{T}\end{cases}$
where $T_{L}$ and $T_{R}$ are strictly increasing, continuous and differentiable functions on $\mathbb{R}^{+}=[0,\infty)$ such that $T_{L}(0)=T_{R}(0)$ and $\mu_{T}\in\mathbb{R}$.
Theorem 2.
Let $(Y,Z)^{\top}$ be a random vector in $\mathbb{R}^{2}$ with joint density $f_{Y,Z}$, marginals densities $f_{Y}$ and $f_{Z}$ and copula density $c$. If the density ratio $f_{Y,Z}(y,z)/(f_{Y}(y)f_{Z}(z))$ can be written in the form
$\frac{f_{Y,Z}(y,z)}{f_{Y}(y)f_{Z}(z)}=g\left(T_{1}(y),T_{2}(z)\right)$
for two v-shaped transformations $T_{1}$ and $T_{2}$ on $\mathbb{R}$ and some function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}$, then we may write $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ on $[0,1]^{2}$ where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are v-transforms given by $\mathcal{V}_{1}(u)=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}(u)$ and $\mathcal{V}_{2}(v)=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}(v)$ and where $c^{*}$ is the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\top}$.
Proof.
The density $c$ is the density of the df of the random variables $(U,V)^{\top}$ where $U=F_{Y}(Y)$ and $V=F_{Z}(Z)$. In McNeil, (2021) it is shown that $F_{T_{1}(Y)}(T_{1}(Y))=\mathcal{V}_{1}(U)$ for the v-transform given by $\mathcal{V}_{1}=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}$; similarly $F_{T_{2}(Z)}(T_{2}(Z))=\mathcal{V}_{2}(V)$ for $\mathcal{V}_{2}=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}$. Hence we may write
$c(u,v)=g\left(T_{1}(F_{Y}^{-1}(u)),T_{2}(F_{Z}^{-1}(v))\right)=g\left(F_{T_{1}(Y)}^{-1}(\mathcal{V}_{1}(u)),F_{T_{2}(Z)}^{-1}(\mathcal{V}_{2}(v))\right).$ (19)
Let $c^{*}$ be the density of the df of $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ and thus the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\prime}$. We can compute $c^{*}$ from (19) using (14). Since $\mathcal{V}_{1}(\mathcal{V}_{1}^{-1}(u))=\mathcal{V}_{1}(\mathcal{V}_{1}^{-1}(u)+u)=u$ and $\mathcal{V}_{2}(\mathcal{V}_{2}^{-1}(v))=\mathcal{V}_{2}(\mathcal{V}_{2}^{-1}(v)+v)=v$ we infer that $c^{*}(u,v)=g\left(F_{T_{1}(Y)}^{-1}(u),F_{T_{2}(Z)}^{-1}(v)\right)$ and hence it follows that $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$. ∎
We give two examples to illustrate the use of Theorem 2.
Example 1 (Symmetric GARCH-type process).
Theorem 2 gives an alternative proof of the fact that the copula density $c_{1}$ in a symmetric GARCH-type process can be written in the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$. The joint density in (6) may be written as
$f_{X_{1},X_{2}}(x,y)=\int_{0}^{\infty}\frac{1}{w\sigma\left(|x|,s\right)}f_{\epsilon}\left(\frac{|y|}{\sigma\left(|x|,s\right)}\right)f_{\epsilon}\left(\frac{|x|}{s}\right)f_{\sigma}(s)\mathrm{d}s$
in terms of the absolute value function, which is a v-shaped transformation on $\mathbb{R}$; moreover the marginal distribution satisfies $f_{X}(x)=f_{X}(|x|)$.
Example 2 (ARCH(1) process with symmetric innovations and leverage).
For the process of ARCH(1) type with leverage and symmetric distribution in the upper left of Figure 3 we observe that the volatility function $\sigma(x)$ is itself a v-shaped transformation with $\mu_{\sigma}=0$, $\sigma_{L}(x)=\sqrt{\alpha_{0}+(\alpha_{1}+\gamma_{1})x^{2}}$ and $\sigma_{R}(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}}$. In this case, it follows from the joint density in (9) that we may write
$\frac{f_{X_{1},X_{2}}(x,y)}{f_{X}(x)f_{X}(y)}=\frac{f_{\epsilon}\left(\frac{|y|}{\sigma(x)}\right)}{\sigma(x)f_{X}(|y|)}=g(\sigma(x),|y|)$
and conclude that the copula density satisfies $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $c_{1}^{*}$ is the copula density of $(\sigma(X_{1}),|X_{2}|)^{\top}$ or, in other words, $(\sigma_{t},|X_{t}|)^{\top}$ for any $t$. To derive $\mathcal{V}_{1}$ we must find the composition of functions $\mathcal{V}_{1}(u)=F_{\sigma}\circ\sigma\circ F_{X}^{-1}(u)$. By using the fact that the distribution function of $\sigma_{t}$ is
$F_{\sigma}(s)=F_{X}\left(\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}}}\right)-F_{X}\left(-\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}+\gamma_{1}}}\right),$
we obtain after some calculation that
$\mathcal{V}_{1}(u)=\begin{cases}F_{X}\left(-\left(1+\frac{\gamma_{1}}{\alpha_{1}}\right)^{\frac{1}{2}}F_{X}^{-1}(u)\right)-u,&u\leqslant 0.5,\\ u-F_{X}\left(-\left(1+\frac{\gamma_{1}}{\alpha_{1}}\right)^{-\frac{1}{2}}F_{X}^{-1}(u)\right),&u>0.5,\end{cases}$
which is an asymmetric function for $\gamma_{1}\neq 0$.
For the classical ARCH(1) process with asymmetric innovation density in the upper right panel of Figure 3 it does not seem to be generally possible to write the copula density in the form $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$. Although the density ratio can be written in the form
$\frac{f_{X_{1},X_{2}}(x,y)}{f_{X}(x)f_{X}(y)}=\frac{f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)}{\sigma(x)f_{X}(y)}=h(\sigma(x),y),$
it does not appear to be possible to find a second v-shaped transformation $T_{2}$ such that $h(\sigma(x),y)=g(\sigma(x),T_{2}(y))$. This conclusion extends to the more general models of first-order GARCH type with asymmetric innovation distributions.
Despite the negative conclusion, copula densities of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ are still promising candidates for approximating or mimicking the implicit copulas of GARCH(1,1)-type models and this is the subject of the next section.
5 GARCH-mimicking copulas
5.1 Mimicking jointly symmetric GARCH copulas
There are a number of ways of finding jointly symmetric copulas that might be capable of mimicking the jointly symmetric copula of a symmetric GARCH process. We could consider:
  1. 1.
    the copula of a spherically symmetric distribution;
  2. 2.
    an equal (50/50) mixture of a radially symmetric copula and its 90-degree rotation;
  3. 3.
    an equal (25/25/25/25) mixture of any copula and its rotations through 90, 180 and 270 degrees;
  4. 4.
    a copula constructed from another copula $C^{*}$ by inverse v-transformation with the symmetric v-transform, i.e. a copula with density of the form $c^{*}(|2u-1|,|2v-1|)$.
In view of Proposition 6, the densities of the copulas arising from the first 3 approaches can also be written in the form $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some implicit copula $C^{*}$ with density $c^{*}$, so the 4th approach should be considered as the overarching approach. Proposition 3, part (ii), and the lower-left panel of Figure 3 give clues about the kind of copula density $c^{*}$ we should be targeting. In particular, we note that there should be strong dependence in the upper right tail and some asymmetry.
An example of (1) is the copula of a spherical t distribution, i.e. a bivariate t distribution located at the origin with identical marginal distributions and correlation equal to zero. We call this copula the spherical t copula and denote it $C^{t}_{\nu}$. Since a spherical t-distributed random vector $(Y,Z)^{\top}$ is jointly symmetric, so is its copula. As the degree-of-freedom parameter tends to zero, this copula becomes increasingly cross-shaped and converges to an equal mixture of the comonotonicity and countermonotonicity copulas (McNeil et al.,, 2022).
It is of interest to identify the copula corresponding to v-transformation of $C^{t}_{\nu}$, in other words, the copula of $(|Y|,|Z|)^{\top}$. We call this copula the absolute spherical t copula (ast) and denote it $C^{|t|}_{\nu}$. By Corollary 1 and (16) its density must satisfy
$c^{|t|}_{\nu}(u,v)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1+v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1+v}{2}\right).$
Integrating any one of these equalities and observing the boundary conditions for a copula gives us an expression for $C^{|t|}_{\nu}$. For example, integrating the final equality gives
$C^{|t|}_{\nu}=4C_{\nu}^{t}\left(\frac{1+u}{2},\frac{1+v}{2}\right)-u-v-1.$ (20)
This one-parameter copula, which appears to be new, interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$). For small degrees of freedom, the density $c^{|t|}_{\nu}$ does bear some resemblance to the lower-left panel of Figure 3, but cannot reproduce the non-exchangeability.
An example of (2) would be an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and the t copula $C^{t}_{\nu,-\rho}$ for some $\rho\geqslant 0$; obviously this contains the spherical t copula $C^{t}_{\nu}$ as the special case where $\rho=0$. Examples of (3) could be constructed using the Gumbel, Clayton or Joe copulas, to take advantage of the strong tail dependence of these copulas in either the upper-right or lower-left corners; this includes the equally-weighted special case of the so-called convex Gumbel copula (Junker and May,, 2005) used by Loaiza-Maya et al., (2018) for financial return data. Mixtures of types (2) and (3) all imply v-transformed copulas $C^{*}$, whose forms may be easily derived using the method we have used for the spherical t copula. If we follow the explicit construction (4), we need to choose a copula $C^{*}$ that has stronger dependence in the upper tail than the lower tail. This suggests considering the Gumbel, Joe or 180-degree-rotated (survival) Clayton copulas.
So far none of these constructions address the issue of non-exchangeability which was clearly evident in certain examples in Figure 2. One possible construction for non-exchangeable copulas is the one proposed by Khoudraji, (1995) and later generalized by Liebscher, (2008). In this construction two copulas $C_{A}$ and $C_{B}$ are combined to give a new copula
$C_{K}(u,v)=C_{A}(u^{1-a_{1}},v^{1-a_{2}})C_{B}(u^{a_{1}},v^{a_{2}})$ (21)
by choosing constants $a_{i}\in[0,1]$ for $i=1,2$ and the resulting copula $C_{K}$ is non-exchangeable when $a_{1}\neq a_{2}$. In our use of this construction we set $C_{B}$ to be the independence copula. We apply the Khoudraji extension to copulas before either mixing them with their rotations or applying inverse v-transformation.
We note, however, that the class of copulas in (21) does have one problematic feature for maximum-likelihood inference; even when $C_{B}$ is the independence copula, in order to evaluate the copula density of $c_{K}$ of $C_{K}$, we need to be able to evaluate, not only the copula density $c_{A}$, but also the copula $C_{A}$ and its partial derivatives. This is a particular challenge for copulas related to the t copula [^3] For example, existing R software for evaluating the distribution function of the t copula imposes some limitations on $\nu$. The algorithms in the R package copula (Hofert et al.,, 2023), which are based on methods by Gentz and Bretz (Gentz and Bretz,, 2009), are limited to integer $\nu$ and this is insufficiently fine-grained for applications of $C^{|t|}_{\nu}$. The C++ library underlying rvinecopulib (Nagler and Vatter,, 2022) imposes the restriction that $\nu\geqslant 2$..
5.2 Simulation study of jointly symmetric case
We explore the ideas of the previous section via a Monte Carlo simulation study. We generate data from the classical symmetric ARCH(1) and GARCH(1,1) processes using the parameter values and innovation distributions summarized in Table 1. We then fit copulas with the property of joint symmetry to the generated realizations of the pairs $(X_{t},X_{t+1})^{\top}$.
In this and all subsequent simulations, the analyses are based on simulated samples of 50000 points. Copulas are estimated by the method of pseudo maximum likelihood (PML) (Genest and Rivest,, 1993); in other words, marginal distributions are estimated empirically, the data are mapped to the unit square using the empirical distribution function and copula parameters are then determined by maximum likelihood. Models are compared according to their AIC values.
The first copula fitted in Table 2 is $C^{t}_{\nu}$, the copula of the spherical t distribution, i.e. the t copula with the correlation parameter fixed at $\rho=0$; the only free parameter is the degree-of-freedom parameter $\nu$. We also label this copula iv-ast, to emphasize that it can be viewed as the inverse v-transformed version of the absolute spherical t copula $C^{|t|}_{\nu}$. The second model is the generalisation of $C^{t}_{\nu}$ consisting of an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and its 90-degree rotation $C^{t}_{\nu,-\rho}$ for $\rho\geqslant 0$.
According to AIC, the $C_{\nu}^{t}$ is the best model in the table in two of six cases, which are the cases where the innovations are Student t4. The mixture of two t copulas is an inferior model to $C_{\nu}^{t}$ in 5 of 6 cases, the only exception being GARCH with normal innovations. The maximum likelihood method for the t mixture returns an estimate $\rho\approx 0$ in those 5 cases, which accounts for the fact that the difference in AIC values is approximately 2.
The next panel of 3 models contains equal mixtures of the Clayton, Gumbel and Joe copulas and their 90, 180 and 270 degree rotations. The Gumbel case is the equally-weighted special case of the convex Gumbel copula of Junker and May, (2005), as also used by Loaiza-Maya et al., (2018). The next 3 copulas in the first panel of Table 2 are inverse-v-transformed versions of the upper tail-dependent Gumbel, Joe and Clayton survival copulas (the latter described as Clayton-180), i.e. copulas with densities of the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of either the Clayton survival copula, the Gumbel copula or the Joe copulas.
Comparing all of the 1-parameter models, the $C_{\nu}^{t}$ (iv-ast) copula is always the best model; second place is usually occupied by the iv-Joe copula, except for ARCH with Gaussian innovations where the iv-Clayton180 copula is favoured. Interestingly, while the iv-Gumbel copula is the worst of the inverse-v-transformed copulas, the mixture of Gumbels is often the best of the mixtures.
The remaining models in Table 2 are non-exchangeable. The fourth panel of the table contains mixture copulas, except that the Khoudraji extension in (21) is applied to each copula before the mixture of the 4 different rotations is constructed. Thus these models introduce two further parameters $a_{1}$ and $a_{2}$. The final panel contains inverse-v-transformed versions of the Khoudraji extensions of the Clayton survival, Gumbel and Joe copulas. It would be appealing to add the Khoudraji construction based on the iv-ast copula, but this model is computationally more demanding and not currently a viable candidate model for practical applications.
The best of the 3-parameter non-exchangeable models are always in the inverse-v-transformed class. The mixture copulas are only really competitive for t2.5 distributed innovations. The best-fitting copulas for Gaussian innovations (iv-K-Clayton-190 for ARCH and iv-K-Joe for GARCH) give estimates of $a_{2}$ which are very close to zero, while the best-fitting copula for t2.5 innovations (iv-K-Joe) give estimates of $a_{1}$ which are very close to zero. This makes sense in relation to Figure 2 since the non-exchangeability in the copulas seems to be caused by a “twist” in different directions.
It is notable that the best model of all is iv-ast in the two cases where the innovations are t4. In the case of ARCH, this is in accord with the near-exchangeability evident in Figures 1 and 2. The estimated degree of freedom parameters are 2.56 for ARCH and 2.67 for GARCH.
Model ARCH(1) GARCH(1,1)
Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
$C^{t}_{\nu}$ (iv-ast) 1 482 0 115 56 0 77
mix-t 2 484 2 117 54 2 79
mix-Clayton 1 1630 426 178 444 284 127
mix-Gumbel 1 867 150 171 181 113 138
mix-Joe 1 1936 593 265 619 424 214
iv-Clayton-180 1 483 146 281 127 98 198
iv-Gumbel 1 1121 444 368 375 300 272
iv-Joe 1 604 50 134 97 58 113
mix-K-Clayton 3 1124 430 91 338 281 52
mix-K-Gumbel 3 728 154 55 152 117 46
mix-K-Joe 3 1634 597 167 585 428 133
iv-K-Clayton-180 3 0 147 145 42 91 101
iv-K-Gumbel 3 552 448 210 283 296 141
iv-K-Joe 3 47 54 0 0 56 0
Table 2: Comparison of copula fits when jointly symmetric copulas are fitted to simulated data from symmetric ARCH(1) and GARCH(1,1) processes by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
5.3 Simulation study of h-symmetric case with leverage
Model ARCH(1) GARCH(1,1)
Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
iv-ast 1 980 224 222 683 262 173
iv-Clayton-180 1 1009 358 363 806 384 299
iv-Joe 1 1114 267 237 727 319 200
iva-ast 2 258 0 131 23 0 68
iva-Clayton-180 2 364 152 280 209 139 202
iva-Joe 2 374 48 152 54 60 100
iv-K-Clayton-180 3 655 361 239 715 386 238
iv-K-Joe 3 672 271 113 619 323 132
iva-K-Clayton-180 4 22 153 131 151 136 112
iva-K-Joe 4 0 52 0 0 61 0
Table 3: Comparison of AIC values when various h-symmetric copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with symmetric innovations and leverage by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
In this section we consider the ARCH(1) and GARCH(1,1) models with symmetric innovations and leverage. According to Proposition 4, the first-order copula $C_{1}$ is h-symmetric but not jointly symmetric. According to Example 2, the ARCH(1) process has a copula density of the form $c(\mathcal{V}_{1}(u),|2v-1|)$ for an asymmetric v-transform $\mathcal{V}_{1}$ with fulcrum at $\delta=0.5$. For the GARCH(1,1) model we don’t have the same theoretical result but we will try using copulas of the same form. The parameterisation of the ARCH(1) model with leverage is the same as in Figure 3 while the parameters for the GARCH(1,1) are given in Table 1.
In Table 3 we include a selection of the copulas from Table 2, omitting the mixture copulas, since the inverse-v-transformed copulas typically do better, and also omitting the iv-Gumbel copula, since iv-Clayton-180 and iv-Joe generally do better. We add copulas with the prefix ‘iva’ (referring to asymmetric inverse v-transform) which are models of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $\mathcal{V}_{1}$ is a member of the family $\mathcal{V}(u;\delta,\kappa)$ in (13) with $\delta=0.5$, so that only one extra unknown parameter $\kappa$ is introduced.
The first panel contains jointly symmetric and exchangeable copulas; the second panel contains h-symmetric copulas based on an exchangeable copula for $c^{*}$ but incorporating an asymmetric v-transform for the first variable; the third panel contains jointly symmetric non-exchangeable copulas using the Khoudraji construction; the final panel contains h-symmetric copulas based on a non-exchangeable $c^{*}$ and incorporating an asymmetric v-transform for the first variable.
For the two models with t4 innovations the best fitting copula is the absolute spherical t copula incorporating an asymmetric v-transform. For the other models the best fitting copula is a Khoudraji-transformed version of the Joe copula incorporating an asymmetric v-transform. Results are very similar for the ARCH(1) and GARCH(1,1) process.
5.4 Simulation study of case of asymmetric innovations
Model ARCH(1) GARCH(1,1)
Copula $p$ leverage leverage
iv-ast 1 168 679 49 436
iv-Clayton-180 1 376 881 208 673
iv-Joe 1 220 728 92 487
ivl-ivl-ast 3 116 201 23 92
ivl-ivl-Clayton-180 3 347 558 189 403
ivl-ivl-Joe 3 164 231 64 140
ivl-ivl-K-Clayton-180 5 336 561 191 403
ivl-ivl-K-Joe 5 168 235 68 144
iv2-iv2-ast 5 0 0 0 0
iv2-iv2-Clayton-180 5 228 293 159 277
iv2-iv2-Joe 5 47 24 45 50
iv2-iv2-K-Clayton-180 7 221 297 160 279
iv2-iv2-K-Joe 7 50 28 49 54
Table 4: Comparison of AIC values when various copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with asymmetric innovations by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
In the final experiment we consider models with skewed t innovations with parameters $\nu=4$ and $\lambda=0.8$. We simulate ARCH(1) and GARCH(1,1) processes, both with and without leverage effects; parameter values are as in Table 1. We have already observed that asymmetric innovations remove all of the previous symmetries we have discussed and so our approach in this section is pure experimentation without the theoretical insights we exploited in the previous experiments.
The first set of 3 models in Table 4 consists of jointly symmetric and exchangeable copulas, included as a benchmark. The iv-ast copula is the best of these but substantial improvements can be made by introducing asymmetry.
The second set of 3 models is based on copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ where $\mathcal{V}(\cdot;\delta)$ is the linear v-transform with fulcrum at $\delta$ (asymmetric when $\delta\neq 0.5$) and $c^{*}$ is either the ast, Clayton survival or Joe copula [^4] It should be noted that full maximum likelihood estimation with respect to all parameters including the parameters of the v-transform presents no problems for these models since the ast, Clayton survival and Joe copula densities do not have asymptotes at $(0,0)$. If this were not the case we can run into a threshold estimation issue, as discussed in Bladt and McNeil, 2022b .. These obviously lower the AIC values considerably at the expense of 2 extra parameters. The third set of 2 models applies the Khoudraji transformation to $c^{*}$ in the case of Clayton and Joe, but this only offers an improvement in one case (Clayton-180 and ARCH(1) process).
The fourth set of models is based on copula densities $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ where $\mathcal{V}(\cdot;\delta,\kappa)$ is the 2-parameter v-transform in (13); such v-transforms are both non-linear and asymmetric. These models give the best results of all, with the model where $c^{*}$ is the ast copula taking first place for all 4 datasets, followed by the model based on Joe. Addition of the Khoudraji transformation in the final set of models again offers no discernible improvement to the models based on Joe. It seems that allowing asymmetric v-transform largely removes the need to consider the Khoudraji transformation, even when the v-transforms are linear.
6 Mimicking GARCH(1,1) processes with simplified d-vines
6.1 Practical issues in the construction of simplified d-vines
In the final sections of this paper we provide some guidance to the construction of simplified d-vine copula models to mimic the behaviour of GARCH-type processes and offer potential alternative models for financial time series. The previous sections have given some insight into the kinds of pair copulas we should consider. We now turn to some of the issues that arise when we want to embed these in simplified vine structures.
Equation (1) in the Introduction implies that the density $c_{(n)}$ of the copula of the variables $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary time series model, such as GARCH, can be written in the form
$c_{(n)}(u_{1},\ldots,u_{n})=\\ \prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\mid F_{X}^{-1}(u_{j-k+1}),\ldots,F_{X}^{-1}(u_{j-1}))$ (22)
where, for $k\in\mathbb{N}$, we define the functions $R_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ and $B_{k}:(0,1)\times(0,1)^{k}$
$R_{k}(x;\bm{u})$ $=$ $F_{X_{k+1}|\bm{X}_{[1:k]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{k}),\ldots,F_{X}^{-1}(u_{1}))$
$B_{k}(x;\bm{u})$ $=$ $F_{X_{1}|\bm{X}_{[2:k+1]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{1}),\ldots,F_{X}^{-1}(u_{k}))$
which we refer to as forward and backward Rosenblatt functions, respectively. The notation is based on the one used in Bladt and McNeil, 2022a with minor changes. Note that, in (22), we adopt the convention that $R_{0}(x;\cdot)=B_{0}(x;\cdot)=x$ for all $x\in(0,1)$.
In practice we want to mimic the copula (22) with a simplified d-vine, i.e. a model where the copula has joint density of the simpler form
$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}\Big{(}B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\Big{)}.$ (23)
In this case the Rosenblatt functions are determined in a recursive, interlacing fashion by $R_{1}(x;u)=h_{1,1}(u,x)$, $B_{1}(x;v)=h_{1,2}(x,v)$ and
[…]
where $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$ are the so-called $h$-functions of the copula $C_{k}$ with density $c_{k}$.
It seems unlikely that we can perfectly mimic a GARCH-type process with a simplified d-vine, since the true copula (22) cannot be shown to conform to the simplifying assumption. However, the weak dependence of $c_{2}(u,v;w)$ on the conditioning variable investigated in Section 3.5 gives some optimism that we could at least approximate a GARCH-type model with a simplified d-vine.
Other than ARCH(1)-type processes, which are Markov processes of order 1, GARCH(1,1)-type processes have infinite memory in the sense that $X_{t}$ depends on all previous values $X_{t-k}$, for $k\in\mathbb{N}$, albeit the serial dependence on past observations dies away relatively quickly; for example, the autocorrelation function of the squared values of the classical GARCH(1,1) process decays like an ARMA(1,1) process. To mimic a GARCH(1,1)-type process, it seems natural to consider d-vine processes with infinite-order partial copula dependence as in Bladt and McNeil, 2022a . These are models which are built from a sequence of pair copulas $(C_{k})_{k\in\mathbb{N}}$ which is parsimoniously parameterized using the Kendall partial autocorrelation function (kpacf) or, in other words, the sequence of Kendall rank correlations of $(C_{k})_{k\in\mathbb{N}}$.
GARCH(1,1) models are very parsimoniously parameterized. For example, a model that incorporates both leverage and skewed Student t innovations would have only 6 parameters ($\alpha_{0}$, $\alpha_{1}$, $\beta_{1}$, $\gamma_{1}$, $\nu$, $\lambda$). In our d-vine analogues we also need to control the number of parameters and the method of parameterization via the kpacf developed in Bladt and McNeil, 2022a can be adapted to our context by applying it to an inverse-v-transformed version of the copula series $(C^{*}_{k})_{k\in\mathbb{N}}$. To be more precise, each pair copula $C_{k}$ used in the sequence defining the d-vine would be constructed to have density $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ where the v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ would be two different, probably asymmetric functions with a fixed form for all $k$; in essence they would take care of the effects caused by leverage and asymmetric innovations in a similar manner to $\gamma_{1}$ and $\lambda$ in the aforementioned GARCH(1,1) model. The copula sequence $(C^{*}_{k})_{k\in\mathbb{N}}$ would contain Joe, Clayton survival or ast copulas with Kendall correlations $\tau_{k}$ that coincided with the terms of the kpacf of a classical Gaussian ARMA(1,1) process, or some other low-order linear process. The combination of ARMA(1,1) and two linear v-transforms would give precisely 4 parameters to describe the d-vine copula process, although two or more parameters would also be required to model the marginal distribution of $(X_{t})_{t\in\mathbb{Z}}$.
We have listed the ast copula $C^{|t|}_{\nu}$ as a candidate pair copula for the sequence $(C^{*}_{k})_{k\in\mathbb{N}}$, but it is not clear whether there is a simple parametric formula to link Kendall’s tau to the degree of freedom parameter $\nu$. Simple calculations from (20) show that
$\tau(\nu)=64\int_{0}^{0.5}\int_{0}^{0.5}C_{\nu}^{t}(u,v)c_{\nu}^{t}(u,v)\mathrm{d}u,\mathrm{d}v-1$ (25)
where $C_{\nu}^{t}$ and $c_{\nu}^{t}$ are the spherical t copula and its density, and that $\tau(\nu)$ ranges in the interval $(0,1)$ as $\nu$ ranges in the interval $(0,\infty)$. An analytical expression for $\tau(\nu)$ would make the ast copula a very attractive choice for these models but, in the absence of such an expression, the Joe copula may be the easier option.
Finally we consider the choice of v-transforms for $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. To estimate a d-vine model by maximum likelihood we have to be able to evaluate a likelihood based on (23) which means evaluating the Rosenblatt forward and backward functions using the recursive expressions (24). This in turn entails evaluating multiple $h$-functions. Moreover, to generate realizations from a copula with joint density (23), we have to be able to invert the Rosenblatt forward functions which entails the evaluation of inverse $h$-functions. To this end it is convenient to have closed form expressions for the copulas $C_{k}$ and their $h$-functions and inverse $h$-functions. For copula densities $c_{k}$ obtained from the inverse-v-transformation construction in (15) this is unfortunately not straightforward, except in the case where we have copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ constructed from two linear v-transforms $\mathcal{V}_{\delta_{1}}$ and $\mathcal{V}_{\delta_{2}}$ and a copula density $c^{*}$ which itself has a tractable joint distribution function and $h$-functions. We investigate this in the next section.
6.2 Tractable processes using linear v-transforms
The following results shows that all the formulas required to work with linear v-transforms are easily obtained.
Proposition 7.
The copula $C$ with density given by $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ is
$C(u,v)=\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}C^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}v+\delta_{2}u-\delta_{1}\delta_{2}$ (26)
where $C^{*}$ is the copula with density $c^{*}$. Writing $h_{1}^{*}$ and $h_{2}^{*}$ for the partial derivatives of $C^{*}$, the corresponding partial derivatives $h_{1}$ and $h_{2}$ of $C$ are
$h_{1}(u,v)$ $=-\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{2}$
$h_{2}(u,v)$ $=-\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}h_{2}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}$
and satisfy the equations
$\mathcal{V}\left(h_{1}(u,v);\delta_{2}\right)$ $=h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ (27)
$\mathcal{V}\left(h_{2}(u,v);\delta_{1}\right)$ $=h_{2}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2})).$
Moreover the inverses $h_{1}^{-1}(u,v)=\{x:h_{1}(u,x)=v\}$ and $h_{2}^{-1}(u,v)=\{y:h_{2}(y,v)=u\}$ are
$h_{1}^{-1}(u,v)$ $=-\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}h_{1}^{*-1}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{2}$
$h_{2}^{-1}(u,v)$ $=-\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}h_{2}^{*-1}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}$
where $h_{1}^{*-1}$ and $h_{2}^{*-1}$ are the inverses of $h_{1}^{*}$ and $h_{2}^{*}$.
Proof.
First note that $\frac{\partial^{2}}{\partial u\partial v}C(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$. We just need to check that the expression (26) defines a copula, which entails checking that $C(0,v)=0$ and $C(1,v)=v$ for any $v$; by symmetry it would follow that $C(u,0)=0$ and $C(u,1)=u$ for any $u$. Now
$C(0,v)=\delta_{1}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v-\delta_{1}\delta_{2}$
and this yields 0 for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$; moreover
$C(1,v)=(\delta_{1}-1)\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v+\delta_{2}(1-\delta_{1})$
and this yields $v$ for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$. The $h$-functions are obtained by differentiation of (26) and the identities (27) are easy to verify. Noting that $h_{1}(u,\delta_{2})=\delta_{2}$ and $h_{2}(u,\delta_{1})=\delta_{1}$, it is also straightforward to derive the inverses of the $h$-functions. ∎
We end by considering the special case of the model with linear v-transforms in which the two v-transforms are identical. We find that the resulting d-vine copula also belongs to the class of vt-d-vines in Bladt and McNeil, 2022b ; these models have been used with real financial data and found to be tractable and effective. The statement of the result and proof use the notation $\mathcal{V}(\bm{u};\delta)$ for the element-by-element v-transformation of a vector $\bm{u}$ with the same linear v-transform.
Theorem 3.
Suppose we construct a d-vine copula (23) using a sequence of copulas $(C_{k})_{k\in\mathbb{N}}$ whose densities are all of the form $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}(u;\delta),\mathcal{V}(v;\delta))$ for a sequence of copulas $(C_{k}^{*})_{k\in\mathbb{N}}$ with densities $(c_{k}^{*})_{k\in\mathbb{N}}$ and a single linear v-transform $\mathcal{V}(\cdot,\delta)$. Then
$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}^{*}\Big{(}B^{*}_{k-1}(\mathcal{V}(u_{j-k};\delta);\mathcal{V}(\bm{u}_{[j-k+1:j-1]};\delta)),R^{*}_{k-1}(\mathcal{V}(u_{j};\delta);\mathcal{V}(\bm{u}_{[j-1:j-k+1]};\delta))\Big{)}$ (28)
where $(R_{k}^{*})_{k\in\mathbb{N}}$ and $(B_{k}^{*})_{k\in\mathbb{N}}$ are the forward and backward Rosenblatt functions corresponding to the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$.
Proof.
The result will follow if we can show that the equalities
$\mathcal{V}\left(B_{k-1}(u;\bm{v});\delta\right)=B^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))\quad\text{and}\quad\mathcal{V}\left(R_{k-1}(u;\bm{v});\delta\right)=R^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$ (29)
hold for any $k\in\mathbb{N}$, $u\in(0,1)$ and $\bm{v}\in(0,1)^{k-1}$. We will argue by induction. The equalities (29) are obviously true for $k=1$ since, by our convention, $B_{0}(u;\cdot)=B_{0}^{*}(u;\cdot)=u$ and $R_{0}(u;\cdot)=R_{0}^{*}(u;\cdot)=u$ for all $u$. Assume the equalities (29) are true when $k=l\in\mathbb{N}$ and consider $k=l+1$. For the first equality we can show, by first using formula (27) in Proposition 7 and then applying the inductive assumption, that
$\mathcal{V}\left(B_{l}(u;\bm{v});\delta\right)$ $=\mathcal{V}\left(h_{l,2}\left(B_{l-1}(u;\bm{v}_{[1,k-1]}),R_{l-1}(v_{k};\bm{v}_{[k-1:1]})\right);\delta\right)$
$=h_{l,2}^{*}\left(\mathcal{V}\left(B_{l-1}(u;\bm{v}_{[1,k-1]});\delta\right),\mathcal{V}\left(R_{l-1}(v_{k};\bm{v}_{[k-1:1]});\delta\right)\right)$
$=h_{l,2}^{*}\left(B^{*}_{l-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v}_{[1,k-1]};\delta)),R^{*}_{l-1}(\mathcal{V}(v_{k};\delta);\mathcal{V}(\bm{v}_{[k-1:1]};\delta))\right)=B^{*}_{l}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$
and the same argument can be applied to the second equality. ∎
Equation (28) describes a vt-d-vine copula process. In such a model the copula density can be evaluated by first applying the v-transform to the entire vector $(u_{1},\ldots,u_{n})$ and then passing the resulting values to the density of a d-vine copula based on the sequence $(C_{k}^{*})_{k\in\mathbb{N}}$. McNeil, (2021) shows that if we apply such a copula process to a series $(X_{t})_{t\in\mathbb{Z}}$ we effectively use the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ to model the serial dependence structure of an element-wise transformation of the time series $(T(X_{t}))_{t\in\mathbb{Z}}$. The case where the copulas are all Gaussian corresponds to the vt-ARMA model in McNeil, (2021), although this tends to be an inferior model to the ones based on non-Gaussian d-vines in Bladt and McNeil, 2022b . The implied canonical form of $T$ for a linear v-transform $\mathcal{V}(\cdot;\delta)$ and a continuous marginal distribution function $F_{X}$ is the v-shaped function on $\mathbb{R}$ given by
$T(x)=\begin{cases}F_{X}^{-1}\left(F_{X}(x)+\mathcal{V}(F_{X}(x);\delta)\right)-\mu&x\leqslant\mu,\\ x-\mu&x>\mu,\end{cases}$ (30)
where $\mu=F_{X}^{-1}(\delta)$ [^5] The function $T$ in (30) is not the unique transformation that gives a time series whose dependence structure may be described by the joint density (28). There is, in fact, an equivalence class of transformations that all lead to the same copula density when paired with $F_{X}$. For example, in the case where $F_{X}$ is symmetric and $\delta=0.5$ we could also take $T(x)=x^{2}$. We pin down the canonical form by fixing the gradient to be 1 for $x>\mu$. See McNeil, (2021) for more details.. For example, in the case where $F_{X}$ is symmetric around the origin and $\delta=0.5$, $\mu=0$ and $T(x)=|x|$. In general $T$ can be thought of as a function that measures the magnitude of a return relative to $\mu$ in a way that can also account for asymmetric effects for returns that are greater or less than $\mu$.
7 Conclusion
This paper has asked two questions. First, what are the bivariate copulas that underlie the serial dependencies and serial conditional dependencies of pairs of variables $(X_{t},X_{t+k})$ in first-order GARCH(1,1) models. Second, can these copulas be mimicked by simpler constructions with a view to developing simplified d-vines that can behave in a similar fashion to GARCH-type models, but offer more flexible solutions for datasets where GARCH works less well.
With regard to the first question we find interesting symmetries in the dependence structures of the models with symmetric innovation distributions, even when leverage effects are introduced, but these symmetries disappear in the presence of asymmetric innovations, which are often necessary in practice to model real-world datasets. What can generally always be said is that the underlying bivariate copulas are cross-shaped, in the sense that there is a tendency for points to accumulate in all fours corners of the unit square, and that this tendency increases at shorter lags as the size of the ARCH effect increases and persists at longer lags as the aggregate size of the ARCH and GARCH effects increases. Heavier-tailed innovation distributions lead to weaker dependence.
However, the true copulas and conditional copulas of GARCH-type processes remain inaccessible for practical use, mainly because the stationary distribution of these models is not available in closed form. This leads us to consider the second question of whether GARCH copulas can be mimicked with simpler copulas. Motivated by the recognition that v-transforms are implicitly embedded in the copulas of symmetric GARCH models, we propose a construction based on v-transforms and positive-dependence copulas with strong dependence in the upper joint tail; these constructions are naturally cross-shaped. Of the well-known copulas, Joe’s copula in combination with asymmetric v-transforms seems the most promising for mimicking GARCH. However, we also identify a new copula - the absolute spherical t copula - which seems to offer an even more promising basis for the construction.
Could these copulas be used to construct simplified d-vines to mimic GARCH processes? To this question we give a more qualified answer. While the Markovian first-order ARCH-type models can definitely be mimicked by a d-vine(1) models, as soon as a GARCH effect is introduced, it is not clear whether the simplifying assumption will restrict our ability to perfectly mimic GARCH behaviour; a first investigation of this question in Section 5 produces some weak evidence that the dependence of the conditional copulas on the conditioning variables may not be particularly strong, but more investigation is required. Moreover, the computational requirements of d-vines, as discussed in Section 6, mean that we may have difficulty in integrating the most promising GARCH-mimicking pair copulas (such as iv2-iv2-ast) into d-vine structures and we may have to compromise and use copulas based on linear v-transforms. Further examination of these issues is required.
Could the GARCH-mimicking copulas be used to construct d-vines to model real financial return data? This is a different question to the question of mimicking GARCH processes. Bladt and McNeil, 2022b  have shown that the class of vt-d-vine processes with linear v-transforms is competitive for modelling many financial return series and, in particular, can give better results for data which depart from the power-lawed marginal behaviour modelled by a GARCH process. The d-vine processes constructed from GARCH-mimicking pair copulas, as described in this paper, can offer additional flexibility because of the use of two v-transforms in their construction, even if we do restrict both to be linear. There is a need for further empirical research to evaluate the potential of such models.
Another issue requiring further research is the construction of d-vines to mimic ARMA-GARCH processes. Real-world financial return series sometimes show serial correlation in the raw data as well as in the squared or absolute value series and this is usually modelled by combining ARMA and GARCH effects. In this paper we have concentrated on pure GARCH-type models and given only a single example in Section 3.4 of an AR-ARCH model. The illustration in the lower-right panel of Figure 3 suggests mixtures of positive dependence copulas and GARCH-mimicking copulas could be successful in addressing combined ARMA and GARCH effects. As a pragmatic alternative, the d-vine models discussed in this paper could be applied to the residuals from an ARMA model instead.